How much money does a typical family spend at McDonald’s per visit ? The amount is a normally distributed random variable with mean $ 17,9 and standard deviation $3,55. What is the amount below which only 10% of families spend at McDonald’s ?


1
Expert's answer
2021-12-16T07:25:37-0500

"P(Z<z)=0.1\to z=-1.28."

"\frac{x-17.9}{3.55}=-1.28."

"z=17.9-1.28*3.55=13.36."

The amount below which only 10% of families spend at McDonald’s is $13.36.